OpenStudy (maddie82502):

what is the solution to the proportion 12/15=8/x A. 9 B. 12 C. 16 D. 10 plz help fan and medal

pooja195 (pooja195):

Cross multiply \[\huge~\rm~\bf \frac{ 12 }{ 15 }=\frac{ 8 }{ x }\] \[\huge~\rm~12x=15 \times 8\] 15 x 8=?

OpenStudy (maddie82502):

120

pooja195 (pooja195):

\[\huge~\rm~12x=120\] divide both sides by 12 x=120/12 120/12=?

OpenStudy (maddie82502):

10

pooja195 (pooja195):

Good so x=10

OpenStudy (maddie82502):

thanks @pooja195

pooja195 (pooja195):

yw ^.^